OpenStudy (anonymous):

Sand is being poured onto the ground from a conveyor belt and forms a conical pile. The sand is poured at a rate of 50 cubic feet per minute. The cone is formed so that its radius is always equal to twice its height. How fast is the height changing at the time the height is 10 feet. Leave answer in terms of pi.

OpenStudy (bahrom7893):

Did I not do this?

OpenStudy (bahrom7893):

dV/dt = 50 (ft^3/min) r = 2h h = 10, r = 20, dh/dt - ?

OpenStudy (bahrom7893):

V = (1/3)pi*r^2*h

OpenStudy (bahrom7893):

r = 2h, so: V = (1/3)pi*(2h)^2*h = (pi/3)*4h^3

OpenStudy (anonymous):

i have like 200 homework questions a week. im not even sure what i have done or not lol

OpenStudy (bahrom7893):

dV/dt = 4*pi*h^2*(dh/dt)

OpenStudy (bahrom7893):

50 = 4*pi*100*(dh/dt) dh/dt = 1/(8*pi)

OpenStudy (bahrom7893):

Oh yea.. it worked apparentl!

OpenStudy (bahrom7893):

*apparently

OpenStudy (anonymous):

thanks for the help bah

OpenStudy (anonymous):

might you know this one too Use differentials to approximate the amount in cubic feet of paint needed to cover a cube of side length 100 feet if thickness of the paint is to be 0.01 feet.